Problem 7-3 Constant Growth Valuation Woidtke Manufacturing's stock currently sells for $18 a share. The stock just paid a dividend of $2.50 a share (i.e., D0 = $2.50), and the dividend is expected to grow forever at a constant rate of 5% a year. What stock price is expected 1 year from now? Round your answer to the nearest cent. $

What is the estimated required rate of return on Woidtke's stock? Round the answer to three decimal places.
Answer #1

First we will calculate estimated required rate of return (k)

We know that,

P0 = D1/ k-g

so,

18 = 2.50*1.05 / k - 5

So k =[ 2.625 / 18] +5%

=19.58%

Now we will compute share price one year from now

P1 = D2 / k - g

=2.625*1.05 / [ 19.58 - 5]

=2.75625/14.58%

=18.90